I have an 77% in my class, if the final is worth 30%, what grade do I have to make on the final to get an 90+% out of the course?
Answer by TimothyLamb(4379)

77(1 - 0.30) + f(0.30) >= 90
77(0.70) + f(0.30) >= 90
53.9 + f(0.30) >= 90
f(0.30) >= 36.1
f >= 36.1/(0.30)
---
f >= 120.333
---
you will not get a class grade of 90%, no matter how well you do on the final.
---
the best class grade you can get is 83.9%, but only if you get 100% on the final.
---
53.9 + 100(0.30) = 83.9
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php